# Viability Constraints on Baryogenesis in $f(R,L_{m},T)$ Gravity

### v1

@@ Abstract
Our study explores gravitational baryogenesis in the context of $f(R,L_{m},T)$ gravity, where $R$ is Ricci scalar, $L_{m}$ represents the Lagrangian density of matter-field and $T$ stands for the metric contraction of $T_{\mu\nu}$. We focus on a linear model: $f(R,L_{m},T)=(\alpha R+\beta L_{m}+\gamma T)$, and examine the parameter constrain for a successful baryon asymmetry generation in four different eras of the cosmos under the consideration of a power-law type cosmic expansion. The computed baryon-to-entropy ratio is found to be consistent with the measured order of asymmetry ratio, $9.42\times 10^{-11}$. In addition, the study is further extended to the generalized framework of gravitational baryogenesis, where the outcome shows strong agreement with the current observational data. Our findings indicate that the $f(R,L_{m},T)$ framework ensures compatible theoretical foundation for producing the observed matter imbalance of the cosmos, thereby emphasizing its potential significance in early–universe cosmology.
@@
preprint: APS/123-QED
@@ I Introduction
One of the biggest enigmas in the modern-day cosmology is why our cosmos is exceedingly made up of matters. This puzzle is broadly referred as baryogenesis [1]. Fundamental theories indicate that the early-universe should have contained matter and antimatter in equal measure, leading to the annihilation of each other [2]. However, the fact that we exist in a matter dominated universe means that something unknown in the early era must have tipped the scales. Analysis of astronomical observations from the large-scale structures (LSS), the Cosmic Microwave Background (CMB) [3] and the predictions of Big Bang Nucleosynthesis (BBN) [4] -collectively suggests a visible cosmos with matter predominance. Antimatter, if present at all, is confined to small and isolated regions, with no evidence for large-scale antimatter domains [5, 6, 7]. This observational reality implies that some physical mechanism must be actively operating to produce the detected matter imbalance in the early universe [8]. In 1967, Sakharov identified three essential criteria that must be fulfilled in order to explain the small excess of matter over antimatter observed in the universe [9]:
@@
    Baryon number violation: In the earliest stages of cosmic evolution, interactions capable of violating baryon number must have taken place. Such processes are necessary to generate a net matter surplus.
@@
    Departure from thermal balance: While in thermal balance, forward and backward reactions progress at identical speed, erasing any imbalance. Hence, at certain epochs, the universe must have undergone through an unbalanced phase of evolution.
@@
Among the various theoretical approaches, gravitational baryogenesis provides a compelling scenario, where the baryonic current, $J^{\mu}$, interacts with the temporal differentiation of $R$, thereby inducing $\mathcal{CP}$-violating effects that establishes the dominance of matter [16]. The coupling term that leads to the violation of $\mathcal{CP}$ is given by:
@@
where, cut-off scale of the effective field theory is $M_{*}$ and $det(g_{\mu\nu})=g$.
@@
The notion of gravitational baryogenesis originates from the work in [16], where dynamical $\mathcal{CP}$- violation drives the observed matter–antimatter imbalance. Based on that idea, Lambiase and Scarpetta [17] showed that $f(R)$ gravity naturally accommodates this mechanism, thereby offering a potential arena for generating the required asymmetry. Subsequently, Ramos at el. [18] addressed the same domain and successfully produced the asymmetry ratio in alignment with observations. Odintsov et al. demonstrated that both Gauss–Bonnet gravity [19] and Loop Quantum Cosmology [20] can successfully reproduce asymmetry ratios consistent with current observational constraint. Other important gravity models include torsion-based $f(T)$ gravity [21] and non-metricity inspired $f(Q,T)$ model [22]. For $f(R,T)$ theory, where the action is generalized to be a function of both the curvature scalar $R$ and the energy–momentum tensor’s trace $T$, studies have shown that the mechanism of gravitational baryogenesis can yield viable baryon-to-entropy ratios consistent with observations [23, 24, 25]. Proposed by Harko at el. [26], this hybrid gravity framework, by incorporating explicit matter–geometry couplings, provides a natural extension of GR to address the baryon asymmetry in the early universe. Within the framework of $f(R,L_{m})$ theory, gravitational baryogenesis has been investigated as a possible explanation for the measured baryon asymmetry, with recent studies indicating that the model is capable of producing baryon-to-entropy ratios compatible with observational bounds in a spatially flat radiation dominated universe [28].
@@
Haghani and Harko [29] proposed the gravity framework $f(R,L_{m},T)$ that combine, expand and generalise two widely studied geometry–matter coupling models $f(R,L_{m})$ and $f(R,T)$ by promoting $R$, $L_{m}$ and $T$ into a single function as $f(R,L_{m},T)$. In this construction, $f(R,T)$, $f(R,L_{m})$ and $f(R)$ are recovered as limiting cases. Arora at el. [30] further studied the stability and energy conditions of the $f(R,Lm,T)$ model. The cosmography of the model under consideration is discussed in [31]. It is important to note that for GR, the asymmetry ratio is in direct proportion to temporal differentiation of $R$. For a flat radiation-driven universe, characterized by the equation of state (EoS) parameter $\omega=\frac{1}{3}$, $\dot{R}$ vanishes and hence Eq. (1) yields no net baryon asymmetry. The aim of this study is to explore $f(R,Lm,T)$ theory as a potential framework for gravitational baryogenesis in a flat universe under a power-law expansion scenario, a setting that has not yet been examined for its implications on matter asymmetry.
@@
The sections are organized in the following way. In Section II, an overview of the $f(R,L_{m},T)$ theory framework is presented along with derivation of the corresponding field equation. Section III & IV is devoted to the formulation of gravitational baryogenesis and generalized gravitational baryogenesis under the setting of modified $f(R,L_{m},T)$ models, where analytical developments are carried out in detail. This section also incorporates graphical plots followed by a critical examination and interpretation of the key findings obtained. Finally, Section V provides the concluding remarks and a summary of the main outcomes of the work.
@@ II.1 Review on $f(R,L_{m},T)$ Gravity
This paper addresses a modified theory of gravity where matter interacts non-minimally with spacetime geometry incorporated through the Lagrangian of matter-field $L_{m}$ and the scalar contraction of $T_{\mu\nu}$, yielding a generalized Lagrangian density expressed as $f(R,L_{m},T)$. Thus, the corresponding action for the non-minimally coupled matter–geometry framework of gravity model can be expressed as [29]:
@@
The dynamics of the cosmological equations are examined for two common choices for the Lagrangian density of matter field: $L_{m}=p$ and $L_{m}=-\rho$. For this work, our analysis is carried out using the former case, $L_{m}=-\rho$.
@@
By varying the action functional (with respect to the metric) to zero, we can deduce the field equation. Thus, the field equation takes the following form:
@@
By scalar contraction of Eq. (5), the trace of the field equation is obtained as:
@@
It’s worth-noting that the Eq. (8) transforms into the Einstein’s Field Equation for GR if we take $f(R,L_{m},T)=f(R)$. Thus, it acts as limiting case of the theory.
@@ II.2 Field equations
The corresponding energy–momentum tensor describing the energy and momentum flux distribution throughout spacetime under the perfect fluid approximation of matter is given by:
@@
By using Eqs. (7) and (8), the modified Friedmann Equations can be derived for the ansatz in Eq. (11) as follows:
@@
It is evident from Eq. (15) that for the case of $L_{m}=T=0$, the $f(R)$ gravity energy conservation condition can be restored.
@@ II.3 Overview of Gravitational Baryogenesis
Cosmological theories of present-time suggests that equivalent amounts of matter and antimatter were created in the primordial universe, leading to a vanishing net baryon number. However, analytical results from astronomical observations [4, 3] and the non-detection of large-scale annihilation radiation [7] strongly imply a supremacy of matter in the visible universe. This matter surplus is commonly characterized through a dimensionless quantity termed as Baryon number-to-entropy ratio (BnER) as follows [36]:
@@
The interaction term in Eq. (1) to induce $\mathcal{CP}$-violation is modified for $f(R,L_{m},T)$ gravity as follows:
@@
The expansion of the cosmos is accompanied by a gradual cooling, which in turn reduces both its temperature and energy density. As the universe cools, it undergoes a sequence of phase evolution, each associated with a distinct balanced state defined by a particular symmetry. These transitions lead to the occurrence of $\mathcal{CP}$-violating interactions that is vital for generating the baryon asymmetry, an outcome with significant consequences for the emergence of elementary particles and the LSS of the cosmos [38, 39, 40].
We assume the cosmic evolution to be described by a scale factor of power-law form [41]:
@@
Adopting a generalised scale factor given by Eq. (22), offers a multifaceted and analytically manageable paradigm for investigating gravitational baryogenesis. Specific choices of $\xi$ recover the standard cosmological epochs, like the the matter-driven era $(\omega=0)$ or radiation-dominated epoch $(\omega=\frac{1}{3})$, as special cases. Concisely, the power-law formulation provides the flexibility to examine baryon asymmetry generation over a wider category of cosmic scenarios, while retaining the generalisation for broader analysis.
@@
$H=\frac{\xi}{t},\hskip 113.81102ptR=\frac{6\xi(2\xi-1)}{t^{2}}.$ (23)
@@ III Gravitational Baryogenesis in $f(R,L_{m},T)$
A particular form of $f(R,L_{m},T)$ gravity needs to be adopted instead of relying on its general expression, in order to solve the Eqs. (13) and (14) analytically. In this work, we focus on a linear model: $f(R,L_{m},T)=\alpha R+\beta L_{m}+\gamma T$ [29, 31, 30]. This analysis aims to investigate whether the model can account for the measured matter imbalance via a $\mathcal{CP}$-breaking coupling. The free parameters of the model under consideration are restricted to a physically viable interval such that the theoretically calculated BnER aligns closely with the BnER inferred from astronomical observations. The methodological approach involves determining $t_{D}$ in terms of $T_{D}$, followed by evaluating the BnER within the constraints of the chosen model using Eq. (20).
III.1 $f(R,L_{m},T)$ Model
For our analysis, we adopt $f(R,L_{m},T)$ model of the following form [29]:
@@ III.1 $f(R,L_{m},T)$ Model
$\frac{\eta_{B}}{s}=\frac{g_{b}\sqrt[4]{g_{*}(\omega+1)(-(\beta+2\gamma+16\pi))T_{D}^{4}}\left(45\sqrt{2}(1-2\xi)\sqrt{g_{*}(\omega+1)(-(\beta+2\gamma+16\pi))T_{D}^{4}}+2\pi\sqrt{15}\sqrt{\alpha}g_{*}(2-3\omega)T_{D}^{4}\right)}{4\ 30^{3/4}\sqrt{\pi}\alpha^{3/4}g_{*}M_{*}^{2}\sqrt{\xi}T_{D}}.$ (27)
As highlighted in [16], the interaction terms of the form given in Eq. (1) naturally arise within the framework of low-energy effective field theory derived from fundamental QFT in curved spacetime, provided that the cut-off scale $(M_{*})$ is close to the reduced Planck mass, $M_{*}=\frac{m_{Pl}}{\sqrt{8\pi}}$, where $(M_{p})$ denotes the Planck mass [23, 24]. The model successfully generates the required matter imbalance ratio if the decoupling temperature $(T_{D})$ coincides with the inflationary scale, $M_{I}$. The inflation scale is estimated to be $M_{I}\simeq 2\times 10^{16}GeV$, inferred from the gravitational-wave signatures reported by LIGO [21, 17, 18]. For the evaluation of Eq. (27), the relevant parameter values are taken as: $M_{*}=2\times 10^{18}GeV$, $T_{D}=2\times 10^{16}GeV$, $g_{*}=106$ and $g_{b}\simeq 1$ [17, 23, 24, 25, 18, 19].
Figure 1: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ for the $f(R,L_{m},T)$ model under consideration in dust regime $(\omega=0)$ for three different $\beta$ choices.
Figure 2: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ for the $f(R,L_{m},T)$ model under consideration in radiation era $(\omega=\frac{1}{3})$ for three different $\beta$ choices.
Figure 3: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ for the $f(R,L_{m},T)$ model under consideration in stiff regime $(\omega=1)$ for three different $\beta$ choices.
Figure 4: Plot of $\frac{\eta_{B}}{s}$ versus $\omega$ for the $f(R,L_{m},T)$ model under consideration in quintessence era $(-1<\omega<\frac{1}{3})$ for three different $\alpha$ choices.
III.1.1 Matter-dominated era ($\omega=0$): Dust case
For the dust case, $\omega=0$, the BnER takes the following form:
$\frac{\eta_{B}}{s}\;\Big{|}_{\omega=0}=\frac{8.41592\times 10^{28}(-\beta-2\gamma-50.24)^{1/4}}{\alpha^{1/4}}-\frac{8.91168\times 10^{-6}(-\beta-2\gamma-50.24)^{3/4}}{\alpha^{3/4}}.$ (28)
Fig. 1 represents the baryon asymmetry ratio $(\frac{\eta_{B}}{s})$ as a function of $\alpha$ for $\beta=-1\times 10^{2}$, $\beta=-5\times 10^{2}$ and $\beta=-1\times 10^{3}$. The value of $\gamma$ is fixed at $\gamma=1$ for matter-driven case. The dashed-horizontal line denotes the measured value of the matter surplus ratio, $\frac{\eta_{B}}{s}=9.42\times 10^{-11}$. It can be concluded from Fig. 1 that the calculated value of BnER aligns closely with observational data for $\alpha$ values roughly in the range ($0.5\times 10^{158},6.0\times 10^{158}$).
III.1.2 Radiation-dominated era ($\omega=\frac{1}{3}$)

### v2

@@ Abstract
Our study explores gravitational baryogenesis in the context of $f(R,L_{m},T)$ gravity, where $R$ is the Ricci scalar, $L_{m}$ represents the Lagrangian density of the matter field and $T$ stands for the metric contraction of $T_{\mu\nu}$. We focus on models: (i) $f(R,L_{m},T)=\alpha R+\beta L_{m}+\gamma T$ and (ii) $f(R,L_{m},T)=\alpha R+\beta L_{m}^{2}+\gamma T^{2}$, to examine the parameter constraints for a successful baryon asymmetry generation in the radiation dominated era of the cosmos under the consideration of a power-law-type cosmic expansion. The computed baryon-to-entropy ratio is found to be consistent with the measured order of the asymmetry ratio $9.42\times 10^{-11}$. In addition, the study is further extended to the generalised framework of gravitational baryogenesis, where the result shows strong agreement with the current observational data. Our findings indicate that the $f(R,L_{m},T)$ framework ensures a compatible theoretical foundation for producing the observed matter imbalance of the cosmos, thus emphasising its potential significance in early–universe cosmology.
@@
Keywords: $f(R,L_{m},T)$ theory, Matter-Antimatter imbalance, Baryon number-to-entropy ratio, Gravitational baryogenesis.
@@ I Introduction
One of the biggest enigmas in the present-day cosmology is why our universe is exceedingly made up of matter compared to antimatter. This puzzle is broadly referred to as baryogenesis [1]. Fundamental theories indicate that the early universe should have contained matter and antimatter in identical measure, leading to the annihilation of each other [2]. However, the fact that we exist in a matter-dominated universe means that something unknown in the early era must have tipped the scales. Analysis of astronomical observations from the large-scale structures (LSS), the Cosmic Microwave Background (CMB) [3], and the predictions of Big Bang Nucleosynthesis (BBN) [4] -collectively suggest a visible cosmos with matter predominance. Antimatter, if present at all, is limited to small and isolated regions, with no evidence for large-scale antimatter domains [5, 6, 7]. This observational reality implies that some physical mechanism must be actively operating to produce the detected matter imbalance in the primordial universe [8]. Sakharov, in 1967, classified three essential criteria that must be met to account for the minute baryonic content imbalance observed in the cosmos [9]:
@@
    Baryon number violation: In the earliest stages of cosmic evolution, interactions capable of violating the baryon number must have taken place. Such processes are necessary to generate a net matter surplus.
@@
    Departure from Thermal Balance: While in thermal balance, forward and backward reactions progress at identical speeds, erasing any imbalance. Hence, at certain epochs, the universe must have undergone an unbalanced phase of evolution.
@@
Among the various theoretical approaches, gravitational baryogenesis provides a compelling scenario, where the baryonic current, $J^{\mu}$, interacts with the derivative of $R$, thereby inducing $\mathcal{CP}$-violating effects that establishes the dominance of matter [16]. The coupling term that leads to the violation of $\mathcal{CP}$ is given by:
@@
where, $M_{*}$ corresponds to the cutoff scale in the effective field framework and $det(g_{\mu\nu})=g$.
@@
The notion of gravitational baryogenesis originates from the work in [16], where dynamical $\mathcal{CP}$- violation drives the observed matter–antimatter imbalance. Based on that idea, Lambiase and Scarpetta [17] showed that $f(R)$ gravity naturally accommodates this mechanism, thereby offering a potential arena for generating the required asymmetry. Subsequently, Ramos at el. [18] addressed the same domain and successfully produced the asymmetry ratio in alignment with observations. Odintsov et al. demonstrated that both Gauss–Bonnet gravity [19, 20] and Loop Quantum Cosmology [21] can successfully reproduce asymmetry ratios consistent with current observational constraints. Other important gravity models include torsion-based $f(T)$ gravity [22] and non-metricity inspired $f(Q,T)$ model [23]. More recently, Samaddar et al. [24] demonstrated baryon asymmetry generation in $f(Q,C)$ gravity. Within the $f(R,T)$ formulation, studies have shown that the procedure of gravitational baryogenesis can yield viable asymmetry ratios consistent with observations [25, 26, 27]. Proposed by Harko at el. [28], this hybrid gravity framework, incorporating explicit geometry-matter couplings, provides an extended formulation of GR to address matter surplus in the primordial cosmos. Under the formalism of $f(R,L_{m})$ gravity, gravitational baryogenesis has been investigated as a possible explanation for the measured baryon asymmetry, with recent studies indicating that the model is capable of producing baryon-to-entropy ratios compatible with observational bounds in a spatially flat radiation dominated universe [29].
@@
Haghani and Harko [30] proposed the foundation of $f(R,L_{m},T)$ gravity that combine, expand and generalise two widely studied geometry–matter coupling models $f(R,T)$ and $f(R,L_{m})$ by promoting $R$, $L_{m}$ and $T$ into a single function. In this construction, $f(R)$, $f(R,T)$ and $f(R,L_{m})$ are recovered as limiting cases. Arora at el. [31] further studied the stability and energy conditions of the $f(R,Lm,T)$ model. The cosmography of the model under consideration is discussed in [32]. It must be noted that for GR, the asymmetry ratio is in direct proportion to temporal differentiation of $R$. For a flat radiation-dominated cosmos, marked by the EoS (equation of state) parameter $\omega=\frac{1}{3}$, $\dot{R}$ vanishes and hence Eq. (1) yields no net baryon asymmetry. The aim of this study is to explore $f(R,Lm,T)$ theory as a potential foundation for gravitational baryogenesis in a flat radiation dominated universe under a power-law expansion scenario, a setting that has not yet been examined for its implications on matter asymmetry.
@@
The sections are organized in the following way. In Section II, an overview of the $f(R,L_{m},T)$ framework is presented along with derivation of the corresponding field equation. Section III & IV is devoted to the formulation of gravitational baryogenesis and the generalized scenario in the realm of modified $f(R,L_{m},T)$ models, where analytical developments are carried out in detail. This section also incorporates graphical plots followed by a critical examination and interpretation of the key findings obtained. Finally, Section V provides the concluding remarks and a summary of the main outcomes of the work.
@@ II.1 Review on $f(R,L_{m},T)$ Gravity
This paper addresses a modified theory of gravity where matter interacts non-minimally with spacetime geometry incorporated through the Lagrangian of matter-field $L_{m}$ and the scalar contraction of $T_{\mu\nu}$, yielding a generalized Lagrangian density articulated as $f(R,L_{m},T)$. Thus, the corresponding action for the non-minimally coupled matter–geometry framework of gravity model can be expressed as [30]:
@@
The dynamics of the cosmological equations are examined for two common alternatives for the Lagrangian density of matter field: $L_{m}=p$ and $L_{m}=-\rho$. For this work, our analysis is carried out using the former case, $L_{m}=-\rho$.
@@
Setting $\delta S_{fRL_{m}T}$ equal to zero, we can deduce the field equation. Thus, the field equation takes the following form:
@@
Scalar contraction of Eq. (5), yields the contracted form of the field equation:
@@
It is worth highlighting that the Eq. (8) transforms into the Field Equation of GR, for the special condition: $f(R,L_{m},T)=f(R)$. Thus, it acts as limiting case of the theory.
@@ II.2 Field equations
The corresponding energy and momentum flux distribution throughout spacetime is described by the energy–momentum tensor (under the perfect fluid approximation of matter) and is given by:
@@
By using Eqs. (7) and (8), the modified Friedmann Equations may be deduced for the ansatz in Eq. (11) as follows:
@@
Eq. (15) clearly shows that for the case: $L_{m}=T=0$, the $f(R)$ gravity energy conservation condition can be restored.
@@ II.3 Overview of Gravitational Baryogenesis
Cosmological theories of present-time suggests that equivalent quantities of matter and antimatter came into existence in the primordial universe, thereby producing no net baryon asymmetry. However, analytical results from astronomical observations [4, 3] and the non-detection of large-scale annihilation radiation [7] strongly imply a supremacy of matter in the visible universe. This matter surplus is commonly characterized through a dimensionless quantity termed as Baryon number-to-entropy ratio (BnER) as follows [37]:
@@
The interaction term in Eq. (1) to induce $\mathcal{CP}$-violation is modified for $f(R,L_{m},T)$ gravity in the following manner:
@@
Because the quantities $R$, $L_{m}$, and $T$ possess different physical dimensions, the coupling constants $\alpha$, $\beta$, and $\gamma$ serve as normalization factors for each term to render them dimensional homogeneity.
@@
The expansion of the cosmos is accompanied by a gradual cooling, which in turn reduces both its temperature and energy density. As the universe cools, it undergoes a sequence of phase evolution, each associated with a distinct balanced state defined by a particular symmetry. These transitions lead to the occurrence of $\mathcal{CP}$-violating interactions that is vital for generating the baryon asymmetry, an outcome with significant consequences for the emergence of elementary particles and the cosmic large-scale structure [39, 40, 41].
We assume the cosmic evolution to be governed by a scale factor of power-law behaviour [42]:
@@
Adopting a generalised scale factor given by Eq. (22), offers a multifaceted and analytically manageable paradigm to investigate gravitational baryogenesis. Specific choices of $\xi$ recover the standard cosmological epochs as special cases. Concisely, the power-law formulation provides the flexibility to examine baryon asymmetry generation over a wider category of cosmic scenarios, while retaining the generalisation for broader analysis.
@@
$H=\frac{\xi}{t},\quad R=\frac{6\xi(2\xi-1)}{t^{2}}.$ (23)
@@ III Gravitational Baryogenesis in $f(R,L_{m},T)$
A particular form of $f(R,L_{m},T)$ gravity needs to be adopted instead of relying on its general expression, in order to solve the Eqs. (13) and (14) analytically. In this work, we focus on models: (i) $f(R,L_{m},T)=\alpha R+\beta L_{m}+\gamma T$ and (ii) $f(R,L_{m},T)=\alpha R+\beta L_{m}^{2}+\gamma T^{2}$ [30, 32, 31]. This analysis aims to investigate whether the model can account for the measured matter imbalance via a $\mathcal{CP}$-breaking coupling. The parameters of the model under consideration are restricted to a physically viable interval such that the theoretically calculated BnER aligns closely with the BnER inferred from astronomical observations. The methodological approach involves determining $t_{D}$ in terms of $T_{D}$, followed by evaluating the BnER within the constraints of the chosen model using Eq. (20).
III.1 Model I
In Model I, the adopted $f(R,L_{m},T)$ model is expressed as follows [30, 32]:
@@ III.1 Model I
$\frac{\eta_{B}}{s}=\frac{g_{b}\sqrt[4]{g_{*}(\omega+1)(-\beta-2\gamma-16\pi)T_{D}^{4}}\left(45\sqrt{2}(1-2\xi)\sqrt{g_{*}(\omega+1)(-\beta-2\gamma-16\pi)T_{D}^{4}}+2\pi\sqrt{15}\sqrt{\alpha}g_{*}(2-3\omega)T_{D}^{4}\right)}{4\times 30^{3/4}\sqrt{\pi}\alpha^{3/4}g_{*}M_{*}^{2}\sqrt{\xi}T_{D}}.$ (27)
As highlighted in [16], the interaction terms of the form given in Eq. (1) naturally arise within the framework of low-energy effective field theory originating from QFT in curved spacetime, given that the cut-off scale $(M_{*})$ is close to the reduced Planck mass, $M_{*}=\frac{m_{Pl}}{\sqrt{8\pi}}$, where $(M_{p})$ denotes the Planck mass [25, 26]. The model successfully generates the required matter imbalance ratio if $T_{D}\leq M_{I}$, $M_{I}$ is the inflationary scale. The inflation scale is estimated to be $M_{I}\simeq 2\times 10^{16}GeV$, inferred from the gravitational-wave signatures reported by LIGO [22, 17, 18]. For the evaluation of Eq. (27), the relevant parameter values are taken as: $M_{*}=2\times 10^{18}GeV$, $T_{D}=2\times 10^{9}GeV$, $g_{*}=106$ and $g_{b}\simeq 1$ [17, 26, 18, 43, 44, 45].
Figure 1: Plot of $\frac{\eta_{B}}{s}$ versus $\alpha$ for the $f(R,L_{m},T)$ model I under consideration in radiation era $(\omega=\frac{1}{3})$ for three different $\beta$ choices.
III.1.1 Radiation-dominated era ($\omega=\frac{1}{3}$)